OpenStudy (anonymous):

Medal&Fan Teresa is making a small table in the shape of a right triangle. If two of the sides will measure about 14 inches each, how many inches should the longest side (c) measure? Use the appropriate formula and round your answer to the nearest tenth. Select one: a. 22.9 in b. 19.8 in c. 16.4 in d. 18.7 in

OpenStudy (johnweldon1993):

A right triangle huh 

OpenStudy (johnweldon1993):

Well you know that using the pythagorean theorem... \[\large c^2 = a^2 + b^2\] So in your case... \[\large c = \sqrt{a^2 + b^2}\] You know that a and b are both 14..so \[\large c = \sqrt{14^2 + 14^2}\] What is your C?

OpenStudy (anonymous):

B ?

OpenStudy (johnweldon1993):

Perfect! :)